# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ 6 Proofs for the Examples in Section 4
where the first identity is a simple exercise and the third identity can be found in e.g. [32, Proposition 1.10.15]. We thus obtain
@@
where the first identity follows from e.g. [32, Corollary 2.7.13] and the third identity can be found in e.g. [32, Proposition 2.6.6]. Moreover, we have $A_{n}(y)=0$ if and only if $\langle y_{1}^{\prime},y\rangle_{F^{\prime},F}=\dots=\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}=0$. In other words, the identity $\ker A_{n}=\ker y_{1}^{\prime}\cap\dots\cap\ker y_{n}^{\prime}$ holds. Combining this with our previous considerations yields
@@
Proof of Lemma 4.3: Let us identify $H$ with $H^{\prime}$ via the canonical and isometric Fréchet-Riesz isomorphism $h\mapsto\hat{h}:=\langle h,\,\cdot\,\rangle_{H}$. Then $(\hat{e}_{n})_{n\geq 1}$ is an ONB of $H^{\prime}$, and hence the assertion follows from Proposition 3.3. ∎
Proof of Proposition 4.4: We first show the existence of a countable $\operatorname{d}_{k}$-dense subset $T_{0}$ of $T$. To this end, we first recall that subsets of separable metric spaces are separable, and therefore, the set
@@
is $\|\cdot\|_{H}$-separable. Consequently, there exists a countable $T_{0}\subset T$ such that
$K_{0}:=\bigl{\{}k(t,\,\cdot\,):t\in T_{0}\bigr{\}}$
is $\|\cdot\|_{H}$-dense in $K_{T}$. To verify, that $T_{0}$ is the desired set, we fix a $t\in T$. Then there exists a sequence $(t_{n})\subset T_{0}$ such that $k(t_{n},\,\cdot\,)\to k(t,\,\cdot\,)$ in $H$. By the definition of $\operatorname{d}_{k}$, this gives $\operatorname{d}_{k}(t,t_{n})\to 0$.
@@
Consequently, every $h\in H$ is $\operatorname{d}_{k}$-continuous. Identifying $H$ with $H^{\prime}$ via the isometric Fréchet-Riesz isomorphism $h\mapsto\hat{h}:=\langle h,\,\cdot\,\rangle_{H}$ we then conclude that the sequence $(\hat{k}(t_{i},\,\cdot\,))$ is separating. Combining Propositions 3.3 and 4.2 then yields the assertion. ∎
@@
see e.g. [11, Theorem V.8.4]. Using the identification (21) of $C(T)^{\prime}$ with ${\cal M}(T)$ and the obvious $\hat{\delta}_{\{t\}}=\delta_{t}$ we conclude that the set $\operatorname{ex}B_{C(T)^{\prime}}$ of extreme points of $B_{C(T)^{\prime}}$ is given by
@@
Krein-Milman’s theorem, see e.g. [32, Corollary 2.10.9], thus gives
@@
where the first identity is trivial and $\operatorname{co}A$ denotes the convex hull of a set $A$. Moreover, $C(T)$ is separable, and hence the relative $w^{*}$-topology on $B_{C(T)^{\prime}}$ is metrizable, see Theorem 2.1. This gives
@@
where we note this identity can also be deduced from [32, Corollary 2.7.13].
@@
where $t_{i,n}\in T$ and $\sum_{i=1}^{m_{n}}|\alpha_{i,n}|\leq 1$. Since $T_{0}\subset T$ is dense, we then find $s_{i,n}\in T_{0}$ with $|t_{i,n}-s_{i,n}|\leq 1/n$ for all $i=1,\dots,m_{n}$ and $n\geq 1$. We define
@@
Let us fix an $f\in C(T)$ and denote its uniform modulus of continuity by $\omega(f,\,\cdot\,)$. Then we have
@@
Then, $\omega\mapsto X_{t}(\omega)$ is measurable for all $t\in T$ since we have a stochastic process. In addition, $t\mapsto X_{t}(\omega)$ is continuous for all $\omega\in\Omega$ since the stochastic process has continuous paths. Moreover, $(T,\operatorname{d})$ is separable, and therefore (68) is ${\cal A}\otimes\mathfrak{B}(T)$-measurable, see e.g. [2, Lemma 4.51]. Fubini-Tonelli’s theorem, see e.g. [27, Theorem 14.16], then shows that the two functions
@@
By Lemma 6.1 we know that $\mathfrak{D}$ is $w^{*}$-dense in $B_{C(T)^{\prime}}$. By (21) and Theorem 5.1 it therefore suffices to show that $\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}$ is Gaussian for all $\hat{\mu}\in C(T)^{\prime}$ of the form $\hat{\mu}=a_{1}\delta_{t_{1}}+\dots+a_{n}\delta_{t_{n}}$, where the coefficients $a_{i}\in\mathbb{R}$ satisfy $|a_{1}|+\dots+|a_{n}|\leq 1$. Let us fix such a $\hat{\mu}$. Defining $Y:=(X_{t_{1}},\dots,X_{t_{n}})$, we then find
@@
Proof of Lemma 6.2: Using (B.3) the first identity follows from
@@
$=\bigl{(}\operatorname{cov}(\delta_{t_{1}}\circ X,Y)\circ(\operatorname{cov}Y)^{\dagger}\circ\operatorname{cov}(Y,X)\bigr{)}(\delta_{t_{2}})$
@@
Proof of Proposition 4.7: Since $T_{\infty}$ is dense in $T$, every $f\in C(T)$ is uniquely determined by its behavior on $T_{\infty}$. Consequently, the sequence $(\delta_{t_{i}})$ is separating. Now the assertion follows from Propositions 3.3 and 4.2
Proof of Theorem 4.8: We already know by Proposition 4.7 that the operators $I_{n}:C(S)\to\mathbb{R}$ defined by
@@
where $I_{S_{n}}:C(T)\to\mathbb{R}^{n}$ is defined by (27). Applying Theorem 3.4 with the spaces $E:=C(T)$ and $F:=C(S)$ gives us a $\mathrm{P}_{Y}$-zero set $N\in\mathfrak{B}(C(S))$ such that identities parts i) to viii) hold true. In particular, we know that $\mathrm{P}_{X|Y}(\,\cdot\,|g)$ is a Gaussian measure on $C(T)$ for all $g\in C(S)\setminus N$.
@@
$\bigl{\langle}\delta_{t_{1}},\bigr{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}=K_{t_{1},S_{n}}K_{S_{n},S_{n}}^{\dagger}K_{S_{n},t_{2}}$
@@
Analogously, part viii) of Theorem 3.4 shows $k_{X|Y}(s,t)=\langle\delta_{t},C_{X|Y}\delta_{s}\rangle_{C(T)^{\prime},C(T)}=0$ for all $t\in T$. ∎
@@ A Appendix: Banach Space Valued Integration
Let $(\Omega,{\cal A})$ be a measurable space, $E$ be a Banach space, and $X:\Omega\to E$. It is well known, that $X$ is strongly measurable if and only if it is weakly measurable and $X(\Omega)$ is separable, see e.g. [10, Theorem E.9] or [21, Theorem 1.1.6]. In addition, if $X$ is strongly measurable, then there exists a sequence $(X_{n})$ of (strongly) measurable step functions $X_{n}:\Omega\to E$ with
@@
for all all $\omega\in\Omega$ and $n\geq 1$, see e.g. [10, Proposition E.2]. Here we note that step functions are also called simple functions. In addition, for step functions all three notions of measurability coincide, and for this reason, we usually speak of measurable step functions. Finally, every pointwise limit $X$ of a sequence $(X_{n})$ of strongly measurable functions is strongly measurable, see e.g. [10, Proposition E.1]. Consequently, we can characterize strong measurability by pointwise approximations with measurable step functions in the sense of (A.1). In particular, the finite sum of strongly measurable functions is strongly measurable and if $f:\Omega\to\mathbb{R}$ is measurable and $X:\Omega\to E$ is strongly measurable, then $f\cdot X:\Omega\to E$ is strongly measurable.
@@
where the integral in the middle is independent of the chosen approximation $(X_{n})$, see e.g. [10, page 399]. Moreover, the limit is, of course, with respect to the norm $\|\cdot\|_{E}$. The set ${{\cal L}_{1}(\mathrm{P},E)}$ of all Bochner integrable $X:\Omega\to E$ is a vector space, see e.g. [10, Proposition E.4], which obviously contains all measurable step functions. Moreover, it is easy to see that
@@
see e.g. [10, Proposition E.5]. As usual, we write ${L_{1}(\mathrm{P},E)}$ for the space of $\mathrm{P}$-equivalence classes, which becomes a Banach space if we modify the norm above in the usual way, see e.g. [10, page 401].
@@
For a proof, we refer to e.g. [10, Proposition E.6], where we note that the strong measurability of $X$ follows from e.g. [10, Proposition E.1].
@@
see e.g. [13, Theorem 6 in Chapter II.2] or [21, Theorem 1.2.4] for a stronger result in this direction that goes back to E. Hille. In particular, for all $X\in{{\cal L}_{1}(\mathrm{P},E)}$, we have both $\langle x^{\prime},X\rangle_{E^{\prime},E}\in{{\cal L}_{1}(\mathrm{P})}$ and the “Pettis property”
@@
This shows that for $X\in{{\cal L}_{1}(\mathrm{P},E)}$ the Bochner integral coincides with the so-called Pettis integral, which is discussed in e.g. [21, Chapter 1.2.c].

### v3

@@ 7 Proofs Related to Filtering Sequences in Section 4
where the first identity almost directly follows from the definition of annihilators, and the third identity can be found in e.g. [34, Proposition 1.10.15]. We thus obtain
@@
where the first identity follows from e.g. [34, Corollary 2.7.13] and the third identity can be found in e.g. [34, Proposition 2.6.6]. Moreover, we have $A_{n}(y)=0$ if and only if $\langle y_{1}^{\prime},y\rangle_{F^{\prime},F}=\dots=\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}=0$. In other words, the identity $\ker A_{n}=\ker y_{1}^{\prime}\cap\dots\cap\ker y_{n}^{\prime}$ holds. Combining this with our previous considerations yields
@@
Proof of Corollary 4.4: Let us identify $H$ with $H^{\prime}$ via the canonical and isometric Fréchet-Riesz isomorphism $h\mapsto\hat{h}:=\langle h,\,\cdot\,\rangle_{H}$. Then $(\hat{e}_{n})_{n\geq 1}$ is an ONB of $H^{\prime}$, and hence the assertion follows from the discussion around (21) and Proposition 4.1. ∎
Proof of Theorem 4.5: We begin by verifying that $(\delta_{t_{i}})\subset F^{\prime}$ is separating. To this end, let us fix an $f\in F$ with $f\neq 0$. Then there exists a $t\in T$ with $f(t)\neq 0$ and then a sequence $(t_{i_{k}})\subset T_{\infty}$ with $\operatorname{d}(t_{i_{k}},t)\to 0$. This gives $f(t_{i_{k}})\to f(t)$ and hence there exists an $i_{k}$ with $\langle\delta_{t_{i_{k}}},f\rangle_{F^{\prime},F}=f(t_{i_{k}})\neq 0$. In other words, $(\delta_{t_{i}})\subset F^{\prime}$ is separating. Applying Proposition 4.3 and then Proposition 4.1 thus gives the assertions. ∎
Proof of Lemma of 4.7: Let us fix an $f\in C^{1}([0,1])$ with $f\neq 0$. Then, if $f$ is constant, we find $y_{0}^{\prime}(f)=f(0)\neq 0$. Moreover, if $f$ is not constant, there exists an $t\in(0,1)$ with $f^{\prime}(t)\neq 0$. Since $f^{\prime}$ is continuous, we then find a $t_{i}$ with $y_{i}^{\prime}(f)=f^{\prime}(t_{i})\neq 0$ as in the proof of Theorem 4.5. ∎
Proof of Corollary 4.8: We first note that the separability can be found in e.g. [42, Lemma 4.33].
Our next goal is to show that each $h\in H$ is continuous. To this end, let $\Phi:T\to H$ be the “canonical feature map” given by $\Phi(t):=k(t,\,\cdot\,)$. Since $k$ is continuous, $\Phi$ is also continuous, see e.g. [42, Lemma 4.29]. For $t\in T$ and $(t_{n})\subset T$ with $t_{n}\to t$, we thus find $h(t_{n})=\langle\Phi(t_{n}),h\rangle_{H}\to\langle\Phi(t),h\rangle_{H}=h(t)$ by the reproducing property of $k$.
Now, applying Theorem 4.5 in combination with the reproducing property gives the assertion. ∎
Proof of Proposition 4.9: We first show the existence of a countable $\operatorname{d}_{k}$-dense subset $T_{\infty}$ of $T$. To this end, we first recall that subsets of separable metric spaces are separable, and therefore, the set
@@
is $\|\cdot\|_{H}$-separable. Consequently, there exists a countable $T_{\infty}\subset T$ such that
$K_{0}:=\bigl{\{}k(t,\,\cdot\,):t\in T_{\infty}\bigr{\}}$
is $\|\cdot\|_{H}$-dense in $K_{T}$. To verify, that $T_{\infty}$ is the desired set, we fix a $t\in T$. Then there exists a sequence $(t_{n})\subset T_{\infty}$ such that $k(t_{n},\,\cdot\,)\to k(t,\,\cdot\,)$ in $H$. By the definition of $\operatorname{d}_{k}$, this gives $\operatorname{d}_{k}(t,t_{n})\to 0$.
@@
Consequently, every $h\in H$ is $\operatorname{d}_{k}$-continuous. ∎
Proof of Theorem 4.10: Since $F$ is separable, Theorem 2.1 shows that there exists a metric $\operatorname{d}_{w^{*}}$ on $B_{F^{\prime}}$ that metrizes the relative $w^{*}$-topology of $B_{F^{\prime}}$. In the following, we fix such a metric $\operatorname{d}_{w^{*}}$. Moreover, we consider the function $m:T\to[0,\infty)$ given by
$m(t):=\begin{cases}\|\delta_{t}\|_{F^{\prime}}^{-1}\,,&\mbox{ if}\delta_{t}\neq 0\,,\\ 0\,,&\mbox{ else. }\end{cases}$
Our first goal is to show that for all $f\in F$ and $t\in T$ we have
$\langle\delta_{t},f\rangle_{F^{\prime},F}=\|\delta_{t}\|_{F^{\prime}}\cdot m(t)\cdot\langle\delta_{t},f\rangle_{F^{\prime},F}\,.$ (72)
Here we note that in the case $\delta_{t}\neq 0$ this identity directly follows from the definition of $m$. Moreover, in the case $\delta_{t}=0$ we have $\langle\delta_{t},f\rangle_{F^{\prime},F}=0$ for all $f\in F$, and therefore both sides of (72) equal 00.
Now, from (72) we conclude that $\|\delta_{t}\|_{F^{\prime}}=\|\delta_{t}\|_{F^{\prime}}\cdot\|m(t)\delta_{t}\|_{F^{\prime}}$, which in turn implies $\|m(t)\delta_{t}\|_{F^{\prime}}\leq 1$ for all $t\in T$. This makes it possible to consider the map
$\Phi:T$ $\to B_{F^{\prime}}\times[0,\infty)$
$t$ $\mapsto\bigl{(}m(t)\delta_{t},\|\delta_{t}\|_{F^{\prime}}\bigr{)}\,.$
In the following, we equip $B_{F^{\prime}}\times[0,\infty)$ with the metric
$\operatorname{d}_{w^{*}}^{+}\bigl{(}y_{1}^{\prime},r_{1}),(y_{2}^{\prime},r_{2})\bigr{)}:=\operatorname{d}_{w^{*}}(y_{1}^{\prime},y_{2}^{\prime})+|r_{1}-r_{2}|\,,\qquad\qquad y_{1}^{\prime},y_{2}^{\prime}\in B_{F^{\prime}},r_{1},r_{2}\in[0,\infty).$
Since $(B_{F^{\prime}},\operatorname{d}_{w^{*}})$ is a separable metric space, we quickly verify that $(B_{F^{\prime}}\times[0,\infty),\operatorname{d}_{w^{*}}^{+})$ is also a separable metric space. Since subsets of separable metric spaces are separable metric spaces, we then see that $(\Phi(T),\operatorname{d}_{w^{*}}^{+})$ is also a separable metric space. Consequently, if we define $\operatorname{d}:T\times T\to[0,\infty)$ by
$\operatorname{d}(s,t):=\operatorname{d}_{w^{*}}^{+}(\Phi(s),\Phi(t))\,,\qquad\qquad s,t\in T,$
then $(T,\operatorname{d})$ becomes a separable pseudo-metric space.
It remains to show that all $f\in F$ are continuous with respect to $\operatorname{d}$. To this end, we fix some $f\in F$ and $t\in T$, as well as a sequence $(t_{n})\subset T$ with $\operatorname{d}(t_{n},t)\to 0$. By the definition of $\operatorname{d}$, $\operatorname{d}_{w^{*}}^{+}$, and $\Phi$ the latter ensures
$\operatorname{d}_{w^{*}}\bigl{(}m(t_{n})\delta_{t_{n}},m(t)\delta_{t}\bigr{)}+\bigl{|}\|\delta_{t_{n}}\|_{F^{\prime}}-\|\delta_{t}\|_{F^{\prime}}\bigr{|}=\operatorname{d}_{w^{*}}^{+}(\Phi(t_{n}),\Phi(t))\to 0\,.$
This in turn ensures $\|\delta_{t_{n}}\|_{F^{\prime}}\to\|\delta_{t}\|_{F^{\prime}}$ and
$\bigl{\langle}m(t_{n})\delta_{t_{n}},f\bigr{\rangle}_{F^{\prime},F}\to\bigl{\langle}m(t)\delta_{t},f\bigr{\rangle}_{F^{\prime},F}\,.$
Combining these two limits with (72) we then find $f(t_{n})=\langle\delta_{t_{n}},f\rangle_{F^{\prime},F}\to\langle\delta_{t},f\rangle_{F^{\prime},F}=f(t)$. ∎
8 Proofs for Section 5
@@
see e.g. [13, Theorem V.8.4]. Using the identification (24) of $C(T)^{\prime}$ with ${\cal M}(T)$ and the obvious $\hat{\delta}_{\{t\}}=\delta_{t}$ we conclude that the set $\operatorname{ex}B_{C(T)^{\prime}}$ of extreme points of $B_{C(T)^{\prime}}$ is given by
@@
Krein-Milman’s theorem, see e.g. [34, Corollary 2.10.9], thus gives
@@
where $\operatorname{co}A$ denotes the convex hull of a set $A$ and the first identity is trivial. Moreover, $C(T)$ is separable, and hence the relative $w^{*}$-topology on $B_{C(T)^{\prime}}$ is metrizable, see Theorem 2.1. This gives
@@
where we note this identity can also be deduced from [34, Corollary 2.7.13].
@@
where $t_{i,n}\in T$ and $\sum_{i=1}^{m_{n}}|\alpha_{i,n}|\leq 1$. Since $T_{0}\subset T$ is dense, we then find $s_{i,n}\in T_{0}$ with $\operatorname{d}(t_{i,n},s_{i,n})\leq 1/n$ for all $i=1,\dots,m_{n}$ and $n\geq 1$. We define
@@
Let us fix an $f\in C(T)$ and denote its uniform modulus of continuity by $\omega(f,\,\cdot\,)$, see e.g. [1, page 232]. Then we have
@@
Then, $\omega\mapsto X_{t}(\omega)$ is measurable for all $t\in T$ since we have a stochastic process. In addition, $t\mapsto X_{t}(\omega)$ is continuous for all $\omega\in\Omega$ since the stochastic process has continuous paths. Combining these observations with the separability of $(T,\operatorname{d})$, we conclude by e.g. [3, Lemma 4.51] that (74) is ${\cal A}\otimes\mathfrak{B}(T)$-measurable. Fubini-Tonelli’s theorem, see e.g. [29, Theorem 14.16], then shows that the two functions
@@
By Lemma 8.1 we know that $\mathfrak{D}$ is $w^{*}$-dense in $B_{C(T)^{\prime}}$. By (24) and Theorem 6.1 it therefore suffices to show that $\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}$ is Gaussian for all $\hat{\mu}\in\mathfrak{D}$. Let us therefore fix a $\hat{\mu}\in\mathfrak{D}$. Then there exist an $n\geq 1$ and $a_{1},\dots,a_{n}\in\mathbb{R}$ with $|a_{1}|+\dots+|a_{n}|\leq 1$ and $\hat{\mu}=a_{1}\delta_{t_{1}}+\dots+a_{n}\delta_{t_{n}}$. Defining $Y:=(X_{t_{1}},\dots,X_{t_{n}})$, we then find
@@
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,\operatorname{d})$ be a compact metric space, and $(X_{t})_{t\in T}$ be a Gaussian processes on $(\Omega,{\cal A},\mathrm{P})$ with continuous paths, mean function $m$, and covariance function $k$. Moreover, let $X:\Omega\to C(T)$ be the Gaussian random variable associated to the process via (26) and $S:=\{s_{1},\dots,s_{n}\}\subset T$. Finally, let
@@
Proof of Lemma 8.2: Note that our notations give $\mathbb{E}Y=(\mathbb{E}X_{s_{1}},\dots,\mathbb{E}X_{s_{n}})=m(S)$. Using (B.3) the first identity then follows from
@@
$=\bigl{(}\operatorname{cov}(\delta_{t_{1}}\circ X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}\bigl{(}\operatorname{cov}(Y,X)(\delta_{t_{2}})\bigr{)}$
@@
Proof of Theorem 5.3: We already know by Corollary 4.6 that the operators $I_{n}:C(S)\to\mathbb{R}$ defined by
@@
where $I_{S_{n}}:C(T)\to\mathbb{R}^{n}$ is defined by (30). Applying Theorem 3.3 with the spaces $E:=C(T)$ and $F:=C(S)$ gives us a $\mathrm{P}_{Y}$-zero set $N\in\mathfrak{B}(C(S))$ such that the identities of parts i) to viii) hold true. In particular, we know that $\mathrm{P}_{X|Y}(\,\cdot\,|g)$ is a Gaussian measure on $C(T)$ for all $g\in C(S)\setminus N$.
@@
$\bigl{\langle}\delta_{t_{1}},\bigr{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}=k(t_{1},t_{2})-K_{t_{1},S_{n}}K_{S_{n},S_{n}}^{\dagger}K_{S_{n},t_{2}}$
@@
Analogously, part viii) of Theorem 3.3 shows
$k_{X|Y}(s,t)=k_{X|Y}(t,s)=\langle\delta_{t},C_{X|Y}\delta_{s}\rangle_{C(T)^{\prime},C(T)}=0$
for all $t\in T$. ∎
@@ A Appendix: Banach Space Valued Integration
Let $(\Omega,{\cal A})$ be a measurable space, $E$ be a Banach space, and $X:\Omega\to E$. It is well known, that $X$ is strongly measurable if and only if it is weakly measurable and $X(\Omega)$ is separable, see e.g. [12, Theorem E.9] or [24, Theorem 1.1.6]. In addition, if $X$ is strongly measurable, then there exists a sequence $(X_{n})$ of (strongly) measurable step functions $X_{n}:\Omega\to E$ with
@@
for all all $\omega\in\Omega$ and $n\geq 1$, see e.g. [12, Proposition E.2]. Here we note that step functions are also called simple functions. In addition, for step functions all three notions of measurability coincide, and for this reason, we usually speak of measurable step functions. Finally, every pointwise limit $X$ of a sequence $(X_{n})$ of strongly measurable functions is strongly measurable, see e.g. [12, Proposition E.1]. Consequently, we can characterize strong measurability by pointwise approximations with measurable step functions in the sense of (A.1). In particular, the finite sum of strongly measurable functions is strongly measurable and if $f:\Omega\to\mathbb{R}$ is measurable and $X:\Omega\to E$ is strongly measurable, then $f\cdot X:\Omega\to E$ is strongly measurable.
@@
where the integral in the middle is independent of the chosen approximation $(X_{n})$, see e.g. [12, page 399]. Moreover, the limit is, of course, with respect to the norm $\|\cdot\|_{E}$. The set ${{\cal L}_{1}(\mathrm{P},E)}$ of all Bochner integrable $X:\Omega\to E$ is a vector space, see e.g. [12, Proposition E.4], which obviously contains all measurable step functions. Moreover, it is easy to see that
@@
see e.g. [12, Proposition E.5]. As usual, we write ${L_{1}(\mathrm{P},E)}$ for the space of $\mathrm{P}$-equivalence classes, which becomes a Banach space if we modify the norm above in the usual way, see e.g. [12, page 401].
@@
For a proof, we refer to e.g. [12, Proposition E.6], where we note that the strong measurability of $X$ follows from e.g. [12, Proposition E.1].
@@
see e.g. [15, Theorem 6 in Chapter II.2] or [24, Theorem 1.2.4] for a stronger result in this direction that goes back to E. Hille. In particular, for all $X\in{{\cal L}_{1}(\mathrm{P},E)}$, we have both $\langle x^{\prime},X\rangle_{E^{\prime},E}\in{{\cal L}_{1}(\mathrm{P})}$ and the “Pettis property”
@@
This shows that for $X\in{{\cal L}_{1}(\mathrm{P},E)}$ the Bochner integral coincides with the so-called Pettis integral, which is discussed in e.g. [24, Chapter 1.2.c].